OpenStudy (anonymous):

how to integrate 1/(0.75-x/200)

OpenStudy (anonymous):

\[\large \int\limits\limits\limits_{ }^{ } \frac{1}{0.75-\frac{v}{200}}\]

OpenStudy (anonymous):

i know the one part would be \[\ln (0.75-\frac{v}{200})\] but how do you do the rest?

OpenStudy (anonymous):

@estudier

OpenStudy (anonymous):

try let \[u= 0.75 -\frac{v}{200}\]

OpenStudy (anonymous):

\[\frac{du}{dx} =-\frac{1}{200} \]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

\[\frac{du}{dx} =-\frac{1}{200} => 200du=-dx=> -200du=dx\] \[\int\limits\frac{1}{u} *-200du => -200\int\limits\frac{1}{u} du\] and np :)